OpenStudy (anonymous):

sqrt 16y^3 - sqrt y^3

OpenStudy (rajee_sam):

\[\sqrt{16y ^{^{3}}} - \sqrt{y ^{3}} = 4\sqrt{y ^{3}} - \sqrt{y ^{3}}\]\[= 3\sqrt{y ^{3}} = 3y \sqrt{y}\]